# Thresholded Lasso for high dimensional variable selection

### v2

@@ 2.1 The thresholding rules
Previously, it has been shown that (11) holds with high probability for the Dantzig selector under the condition of a Uniform Uncertainty Principle (UUP), where the UUP states that for all $s$-sparse sets $J$, the columns of $X$ corresponding to $J$ are almost orthogonal [9]. Then under the settings of Lemma 2.7, we have $\theta_{|I|,|{S_{{\mathcal{D}}}}|}\leq\delta_{2s}<1$, where $|I|+|S_{{\mathcal{D}}}|\leq 2s$; cf. (31). The tight analysis in Theorem 2.1 for the Thresholded Lasso estimator is motivated by the sparse oracle inequalities on the Gauss-Dantzig selector under the UUP, which is originally shown to hold in a conference paper by [47].
2.2 Discussions
The sparse eigenvalue conditions in the present work are considerably relaxed from the incoherence condition (UUP) in Definition I.1: The UUP condition ensures that the $\textnormal{{RE}}(s,1,X)$ condition as in (3) holds with
@@ 2.2 Discussions
cf. [3]. Besides the sparse eigenvalue conditions, Theorem 2.1 requires $\textnormal{{RE}}(s_{0},4,X)$ to be satisfied, which depends on the essential sparsity $s_{0}$ rather than $s=\left\lvert\mathop{\text{\rm supp}\kern 0.86108pt}(\beta)\right\rvert$; cf. Section 3 for detailed discussions. In the supplementary Theorem I.3, we show the corresponding result for the Gauss-Dantzig selector under the UUP for completeness.
In summary, Lemmas 2.3 and 2.7 ensure that the general thresholding rules with threshold level at about $\lambda\sigma$ achieve the following property: although we cannot guarantee the presence of variables indexed by $S_{R}=\{j:|\beta_{j}|<\sigma\sqrt{\log p/n}\}$ to be included due to their lack of strength, we will include in $I$ most variables in $S\setminus S_{R}$ such that the OLS estimator based on model $I$ achieves the oracle bound (11). This goal is accomplished despite some variables from the support set $S$ are missing from the model $I$, since their overall $\ell_{2}$-norm $\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}$ is bounded in (21). As mentioned, Proposition 4.1 (by setting $c^{\prime}=1$) shows that the number of variables in $\beta$ that are larger than or equal to $\sqrt{\log p/n}\sigma$ in magnitude is bounded by $2s_{0}$. In hindsight, it is clear that we wish to retain most of them by keeping $2s_{0}$ variables in the model $I$. Indeed, suppose $D_{1}s_{0}<(c_{0}s_{0})\wedge(s-s_{0})$. Then, by choosing $t_{0}\asymp\lambda\sigma$ on event ${\mathcal{T}}_{a}$, we are guaranteed to obtain under the settings of Theorem 2.1 (and Lemma 2.3),
@@ 2.3 Background and related work
Prior to our work, a similar two-step procedure, namely, the Gauss-Dantzig selector, has been proposed and empirically studied by [9]. This paper builds upon the methodology originally developed in a conference paper by the present author [47].  [47] obtains oracle bounds in the spirit of Theorem 2.1 for the Gauss-Dantzig selector, under the stronger restricted isometry type of conditions as originally proposed by [9]; cf. the supplementary Theorem I.3.
@@
Subsequent development. This choice of the threshold parameter identified in [47, 48] and the current paper has deep connection with the classic and current literature on model selection [16, 14, 41, 42, 39].  [42] proves the minimax concave penalty (MCP) procedure is selection consistent under a sparse Riesz condition and an information requirement in the sense of (4). Nonconvexity of the minimization problem cause computational and analytical difficulties; Compared to the elegant yet complex method in MCP, the Thresholded Lasso procedures [47, 48] provide a much simpler framework, which is desirable from the practical point of view, with overall good performance. This is confirmed in a subsequent study by [40].
@@ 3 Proof sketch for the main result
We will now describe the main ideas of our analysis in this section. Combining Theorem 2.4 with Lemmas 2.5 and 2.3 allows us to prove Theorem 2.1, which we will elaborate in more details in Section A.1. For now, we highlight the important differences between our results and a previous result on the Lasso by [3] (cf. Theorem 7.2), which we refer to as the BRT results. While a bound of $O_{P}(\lambda\sigma\sqrt{s})$ on the $\ell_{2}$ loss as obtained by [3] makes sense when all signals are strong, significant improvements are needed for the general case where $\beta_{\min}$ is not bounded from below.
@@
(b) We note that the lower sparse eigenvalue condition $\Lambda_{\min}(2s)>0$ (6) is implied by, and hence is weaker than the $\textnormal{{RE}}(s,3,X)$ condition. Moreover, it is possible to prove Theorem 2.1 even if we leave condition (6) out. In particular, we note that so long as $|I|\leq 2s_{0}$, then $\textnormal{{RE}}(s_{0},4,X)$ already implies that (34) holds for $c=2$ [3]; cf. Remark 2.9. We will not pursue such optimizations in the present work, as in general, our goal is to bound the model size: $|I|\leq cs_{0}$ for some constant $c>0$ which need not to be upper bounded by 222.
@@
(c) We note that in the $\textnormal{{RE}}(s,3,X)$ condition as required by [3] to achieve the $\ell_{2}$ loss of $O_{P}(\lambda\sigma\sqrt{s})$: while $k_{0}=3$ is chosen, they fix sparsity at $s$ instead of $s_{0}$, which is not ideal when $s_{0}$ is much smaller than $s$. We emphasize that in the $\textnormal{{RE}}(s_{0},4,X)$ condition that we impose, $k_{0}=4$ is rather arbitrarily chosen; in principle, it can be replaced by any number that is strictly larger than 333. In the context of compressed sensing, RE conditions can also be taken as a way to guarantee recovery for anisotropic measurements [31]. Results by [31] reveal that for RE conditions with a smaller $s_{0}$, we need correspondingly smaller sample size $n$ in order for the random design matrix $X$ of dimension $n\times p$ to satisfy such a condition, when the independent row vectors of $X_{i},i=1,\ldots,n$ have covariance $\Sigma(X_{i})={\textbf{E}}X_{i}\otimes X_{i}={\textbf{E}}X_{i}X_{i}^{T}$ satisfying $\textnormal{{RE}}(s_{0},(1+\varepsilon)k_{0},\Sigma^{1/2})$ condition in the sense that (37) holds for any $\varepsilon>0$, for all $\upsilon\not=0$,
@@
$\frac{1}{K(s_{0},(1+\varepsilon)k_{0},\Sigma^{1/2})}\stackrel{{\scriptstyle\triangle}}{{=}}\min_{\stackrel{{\scriptstyle J_{0}\subseteq[p]}}{{|J_{0}|\leq s_{0}}}}\min_{\left\lVert\upsilon_{J_{0}^{c}}\right\rVert_{1}\leq k_{0}\left\lVert\upsilon_{J_{0}}\right\rVert_{1}}\;\;\frac{\left\lVert\Sigma^{1/2}\upsilon\right\rVert_{2}}{\left\lVert\upsilon_{J_{0}}\right\rVert_{2}}>0.$ (37)
@@
(d) We impose an explicit upper bound on $\Lambda_{\max}(2s)$, which is absent from the paper by [3], in order to obtain the tighter bounds in the present work for both the Lasso and the Thresholded Lasso. This condition is required by our OLS refitting procedure as stated in Lemmas 2.7 and 2.8. This is consistent with the fact that known oracle inequalities for the Dantzig and Gauss-Dantzig selectors are proved under the UUP which impose tighter upper and lower sparse eigenvalue bounds in the sense that $\theta_{s,2s}+\delta_{2s}<1$.
3.1 Proof sketch of Theorem 2.4
@@
We specify parameters $\lambda_{n}$ and $t_{0}$ in Theorem 2.4 and Lemma 2.3 respectively. We now sketch a proof of Theorem 2.4, where we elaborate on the $\ell_{p},p=1,2$, loss on $h_{T_{0}}$ and $h_{T_{0}^{c}}$, and their implications on variable selection, where recall $h=\beta_{\text{\rm init}}-\beta_{T_{0}}$ and $T_{0}$ denotes the largest $s_{0}$ coordinates of $\beta$ in absolute values. Improving the bounds on each component will result in a tighter upper bound on controlling the bias. Specifically, these bounds ensure that under the RE and sparse eigenvalue conditions, both the $\ell_{2}$ loss on the set $T_{0}$ of significant coefficients and the $\ell_{1}$ (and the $\ell_{2}$) norm of the estimated coefficients on $T_{0}^{c}$ are tightly bounded with respect to $\left\lvert T_{0}\right\rvert$ for the Lasso estimator: hence achieving an oracle inequality on the $\ell_{2}$ loss in the sense of (11). Specifically, we will show in the supplementary Section H that for some constants $D^{\prime}_{0}$ as in (2.1) and $D_{1}$ in (26), on ${\mathcal{T}}_{a}$,
$\left\lVert h_{T_{0}}\right\rVert_{2}=\left\lVert(\beta_{\text{\rm init}}-\beta)_{T_{0}}\right\rVert_{2}\leq D^{\prime}_{0}\lambda\sigma\sqrt{s_{0}}\;\text{ and }\;\left\lVert h_{T_{0}^{c}}\right\rVert_{1}=\left\lVert\beta_{\text{\rm init},T_{0}^{c}}\right\rVert_{1}\leq D_{1}\lambda\sigma s_{0}.$
@@
First, by definition of $h$, $T_{0}$, we aim to keep variables in $T_{0}$, while for variables outside of $T_{0}^{c}$, we may need to trim these off. It is also clear by Lemma 2.3 that we cannot cut too many “significant” variables in $T_{0}$ by following the thresholding rules in our proposal; for example, for those that are $\geq\lambda\sigma\sqrt{s_{0}}$, we can cut at most a constant number of them. Let $T_{1}$ denote the $s_{0}$ largest positions of $h$ in absolute values outside of $T_{0}$. So what do we do with those in $T_{1}$? The fate of these variables is pretty much up to the choice of the threshold for a given $\beta_{\text{\rm init}}$, knowing these are the largest in magnitude in $h$ (and $\beta_{\text{\rm init}}$) outside of $T_{0}$ and hence most likely to be included in model $I$. Moreover, even if we were able to retain all variables in $T_{0}$, we will include at least some variables in $T_{1}$ when the selection set has size $\left\lvert I\right\rvert>s_{0}$; in fact, we will include nearly all of $T_{1}$ when $I$ has close to $2s_{0}$ variables. Then we have the following bounds on selections from variables in $T_{0}^{c}$:
@@
$\left|I\cap T^{c}_{0}\right|$$\leq$ ${\left\lVert h_{T_{0}^{c}}\right\rVert_{1}}/{t_{0}}\leq D_{1}\lambda_{n}s_{0}/(f_{0}\lambda_{n})\leq(D_{1}/f_{0})s_{0}\quad\text{in case}\quad t_{0}=f_{0}\lambda_{n},$
@@
for some $f_{0}>0$. Before we continue, we state Lemma 3.1, which is the same (up to normalization) as Lemma 3.1 [9], to illuminate the roles of sets $T_{0},T_{1}$ in the overall bounds on $\left\lVert h\right\rVert_{2}$. We note that in their original statement, the UUP condition is assumed; a careful examination of their proof shows that it is a sufficient but not necessary condition; indeed we only need to assume that sparse eigenvalues are bounded, namely, $\Lambda_{\min}(2s_{0})>0$ and $\Lambda_{\max}(2s_{0})<\infty$. Moreover, we state Lemma 3.2.
@@
Let $h=\beta_{\text{\rm init}}-\beta_{T_{0}}$, where $T_{0}$ denotes locations of the $s_{0}$ largest coefficients of $\beta$ in absolute values. Here $\beta_{T_{0}}$ is the restriction of $\beta$ to the set $T_{0}$. Let $T_{1}$ denote the $s_{0}$ largest positions of $h$ in absolute values outside of $T_{0}$. Let $T_{01}:=T_{0}\cup T_{1}$. Suppose $\Lambda_{\min}(2s_{0})>0$ and $\Lambda_{\max}(2s_{0})<\infty$. Then
$\left\lVert h_{T_{01}}\right\rVert_{2}$$\leq$ $\frac{1}{\sqrt{\Lambda_{\min}(2s_{0})}}\left\lVert Xh\right\rVert_{2}/\sqrt{n}+\frac{\theta_{s_{0},2s_{0}}}{\Lambda_{\min}(2s_{0})}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/\sqrt{s_{0}},$
@@
${\mathcal{C}}(s_{0},k_{0}):=\left\{x\in\mathbb{R}^{p},\exists{J\subseteq[p],|J|=s_{0}}\text{ s.t.}\left\lVert x_{J^{c}}\right\rVert_{1}\leq k_{0}\left\lVert x_{J}\right\rVert_{1}\right\}$ (39)
@@
For $u=(u_{1},\ldots,u_{n})\in\mathbb{R}^{n}$, define the empirical norm of $u$ by
$\left\lVert u\right\rVert_{n}^{2}=\left\lVert u\right\rVert_{2}^{2}/n.$
The proof of the original Theorem 5.1 in [48] draws upon techniques from a concurrent work by [36] and uses an elementary inequality (H.1) for the Lasso. Denote by $\delta=\beta_{\text{\rm init}}-\beta=:\hat{\beta}-\beta$. Similar to the proof in [48], we have a deterministic proof on event ${\mathcal{T}}_{a}$, except that now on ${\mathcal{T}}_{a}$,
$\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}+\lambda_{n}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\leq\left\lVert X\beta_{T_{0}^{c}}\right\rVert^{2}_{n}+3\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}.$ (40)
The full proof of Theorem 2.4 appears in the supplementary Section H. The current proof replaces (H.1) for the Lasso with the following updated inequality (126) from [11], cf. Eq(20) therein, where we set $\bar{\beta}=\beta_{0}:=\beta_{T_{0}}$ and $\bar{\delta}:=\hat{\beta}-\bar{\beta}=\hat{\beta}-\beta_{0}:=h$,
$+\frac{2}{n}\epsilon^{T}X(\hat{\beta}-\beta_{0})+2\lambda_{n}(\left\lVert\beta_{0}\right\rVert_{1}-\lVert\hat{\beta}\rVert_{1})$
Now we differentiate between three cases under event ${\mathcal{T}}_{a}$.
@@
    $\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}\geq\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}.$ (42)
    Then
@@
    Moreover, the bounds on $\left\lVert X\delta\right\rVert_{n}$ and $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ ( $\left\lVert h_{T_{0}}\right\rVert$) follow from (45):
@@
    where $\left\lVert X\beta_{T_{0}^{c}}\right\rVert_{n}\leq\lambda\sigma\sqrt{s_{0}\Lambda_{\max}(s-s_{0})}$. Then we have
@@
    $\left\lVert h_{T_{0}}\right\rVert_{2}\leq K(s_{0},4)\left\lVert Xh\right\rVert_{n}\leq\lambda_{n}\sqrt{s_{0}}K(s_{0},4)^{2},\;\text{ and }\;$ (49)
@@
[…]
@@
Using (45), (137), and (51), we conclude that (22) holds. Lemma 3.4 provides the upper bound on $\left\lVert h_{T_{01}}\right\rVert_{2}$, leading to the expression of $D_{0}$ as in (2.1); Moreover, combining (46), (44), (50), and (52), we have the expression of (26) for $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$, namely, $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\leq D_{1}\lambda\sigma s_{0}$.
Next, we state Lemma 3.3. We combine the upper bounds in Lemmas 3.1 and 3.3 in the proof of Lemma 3.4. This allows us to bound $\left\lVert h\right\rVert_{2}$ in view of Lemma 3.1.

### v3

@@ 2.2 Discussions
Then under the settings of Lemma 2.7, we have $\theta_{|I|,|{S_{{\mathcal{D}}}}|}\leq\delta_{2s}<1$, where $|I|+|S_{{\mathcal{D}}}|\leq 2s$; cf. (32). The sparse eigenvalue conditions in the present work are considerably relaxed from the incoherence condition (UUP) in Definition G.1: The UUP condition ensures that the $\textnormal{{RE}}(s,1,X)$ condition as in (3) holds with
@@
cf. [3]. Besides the sparse eigenvalue conditions, Theorem 2.1 requires $\textnormal{{RE}}(s_{0},4,X)$ to be satisfied, which depends on the essential sparsity $s_{0}$ rather than $s=\left\lvert\mathop{\text{\rm supp}\kern 0.86108pt}(\beta)\right\rvert$; cf. Section 3 for detailed discussions. In the supplementary Theorem G.3, we show the corresponding result for the Gauss-Dantzig selector under the UUP for completeness.
In summary, Lemmas 2.3 and 2.7 ensure that the general thresholding rules with threshold level at about $\lambda\sigma$ achieve the following property: although we cannot guarantee the presence of variables indexed by $S_{R}=\{j:|\beta_{j}|<\sigma\sqrt{\log p/n}\}$ to be included due to their lack of strength, we will include in $I$ most variables in $S\setminus S_{R}$ such that the OLS estimator based on model $I$ achieves the oracle bound (11). This goal is accomplished despite some variables from the support set $S$ are missing from the model $I$, since their overall $\ell_{2}$-norm $\left\lVert\beta_{{\mathcal{D}}}\right\rVert_{2}$ is bounded in (21).
As mentioned, Proposition 4.1 (by setting $c^{\prime}=1$) shows that the number of variables in $\beta$ that are larger than or equal to $\sqrt{\log p/n}\sigma$ in magnitude is bounded by $2s_{0}$. In hindsight, it is clear that we wish to retain most of them by keeping $2s_{0}$ variables in the model $I$. Indeed, suppose $D_{1}s_{0}<(c_{0}s_{0})\wedge(s-s_{0})$. Then, by choosing $t_{0}\asymp\lambda\sigma$ on event ${\mathcal{T}}_{a}$, we are guaranteed to obtain under the settings of Theorem 2.1 (and Lemma 2.3),
@@ 2.3 Background and related work
Prior to our work, a similar two-step procedure, namely, the Gauss-Dantzig selector, has been proposed and empirically studied by [9]. This paper builds upon the methodology originally developed in a conference paper by the present author [47]. Inspired by [9][47] obtains oracle bounds in the spirit of Theorem 2.1 for the Gauss-Dantzig selector, under the stronger restricted isometry type of conditions as originally proposed by [9]; cf. the supplementary Theorem G.3.
@@
Subsequent development. This choice of the threshold parameter identified in [47, 48] and the current paper has deep connection with the classic and current literature on model selection [16, 14, 41, 42, 39].  [42] proves the minimax concave penalty (MCP) procedure is selection consistent under a sparse Riesz condition and an information requirement in the sense of (4). Nonconvexity of the minimization problem causes computational and analytical difficulties; Compared to the elegant yet complex method in MCP, the Thresholded Lasso procedures [47, 48] provide a much simpler framework, which is desirable from the practical point of view, with overall good performance. This is confirmed in a subsequent study by [40].
@@ 3 Proof sketch for the main result
We will now describe the main ideas of our analysis in this section. Combining Theorem 2.4 with Lemmas 2.5 and 2.3 allows us to prove Theorem 2.1, which we will elaborate in more details in Section 6.1. For now, we highlight the important differences between our results and a previous result on the Lasso by [3] (cf. Theorem 7.2), which we refer to as the BRT results. While a bound of $O_{P}(\lambda\sigma\sqrt{s})$ on the $\ell_{2}$ loss as obtained by [3] makes sense when all signals are strong, significant improvements are needed for the general case where $\beta_{\min}$ is not bounded from below.
@@
(b) We note that the lower sparse eigenvalue condition $\Lambda_{\min}(2s)>0$ (6) is implied by, and hence is weaker than the $\textnormal{{RE}}(s,3,X)$ condition. Moreover, it is possible to prove Theorem 2.1 even if we leave condition (6) out. In particular, we note that so long as $|I|\leq 2s_{0}$, then $\textnormal{{RE}}(s_{0},4,X)$ already implies that (35) holds for $c=2$ [3]; cf. Remark 2.9. We will not pursue such optimizations in the present work.
@@
(c) We note that in the $\textnormal{{RE}}(s,3,X)$ condition as required by [3] to achieve the $\ell_{2}$ loss of $O_{P}(\lambda\sigma\sqrt{s})$: while $k_{0}=3$ is chosen, they fix sparsity at $s$ instead of $s_{0}$, which is not ideal when $s_{0}$ is much smaller than $s$. We emphasize that in the $\textnormal{{RE}}(s_{0},4,X)$ condition that we impose, $k_{0}=4$ is rather arbitrarily chosen; in principle, it can be replaced by any number that is strictly larger than $3$. In the context of compressed sensing, RE conditions can also be taken as a way to guarantee recovery for anisotropic measurements [31]. Results by [31] reveal that for RE conditions with a smaller $s_{0}$, we need correspondingly smaller sample size $n$ in order for the random design matrix $X$ of dimension $n\times p$ to satisfy such a condition, when the independent row vectors of $X_{i},i=1,\ldots,n$ have covariance $\Sigma(X_{i})={\textbf{E}}X_{i}\otimes X_{i}={\textbf{E}}X_{i}X_{i}^{T}$ satisfying $\textnormal{{RE}}(s_{0},(1+\varepsilon)k_{0},\Sigma^{1/2})$ condition in the sense that for any $\varepsilon>0$, for all $\upsilon\not=0$,
@@
$\frac{1}{K(s_{0},(1+\varepsilon)k_{0},\Sigma^{1/2})}\stackrel{{\scriptstyle\triangle}}{{=}}\min_{\stackrel{{\scriptstyle J_{0}\subseteq[p]}}{{|J_{0}|\leq s_{0}}}}\min_{\left\lVert\upsilon_{J_{0}^{c}}\right\rVert_{1}\leq k_{0}\left\lVert\upsilon_{J_{0}}\right\rVert_{1}}\;\;{\left\lVert\Sigma^{1/2}\upsilon\right\rVert_{2}}/{\left\lVert\upsilon_{J_{0}}\right\rVert_{2}}>0.$
@@
(d) We impose an explicit upper bound on $\Lambda_{\max}(2s)$, which is absent from the paper by [3], in order to obtain the tighter bounds in the present work for both the Lasso and the Thresholded Lasso. This condition is required by our OLS refitting procedure as stated in Lemmas 2.7 and 2.8. This is consistent with the fact that known oracle inequalities for the Dantzig and Gauss-Dantzig selectors are proved under the UUP which impose tighter upper and lower sparse eigenvalue bounds in the sense that $\theta_{s,2s}+\delta_{2s}<1$. See also [43] and the exposition in Section 4.3.
3.1 Proof sketch of Theorems 2.1 and 2.4
@@
We specify the initial Lasso estimator $\beta_{\text{\rm init}}$, and the parameters $\lambda_{n}$, $t_{0}$ in Theorem 2.4 and Lemma 2.3 respectively. Throughout this section, let $h=\beta_{\text{\rm init}}-\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$ be as in Lemma 2.3, where $T_{0}$ denotes the largest $s_{0}$ coordinates of $\beta$ in absolute values. Let $T_{1}$ be the $s_{0}$ largest positions of $h$ outside of $T_{0}$. Denote by $T_{01}=T_{0}\cup T_{1}$. In this section, we elaborate on the $\ell_{p},p=1,2$, loss on $h_{T_{0}}$ and $h_{T_{0}^{c}}$, and their implications on variable selection using thresholding. As mentioned, improving the bounds on each component of $h$ will result in a tighter upper bound on controlling the bias; cf. Lemma 2.3. Specifically, we will show in the supplementary Section D that on ${\mathcal{T}}_{a}$, (19) holds for $h=\beta_{\text{\rm init}}-\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$, where $D^{\prime}_{0}$ and $D_{1}$ are as in (26) and (27).
These bounds ensure that under the RE and sparse eigenvalue conditions, both the $\ell_{2}$ loss on the set $T_{0}$ of significant coefficients, namely, $\left\lVert h_{T_{0}}\right\rVert_{2}=\left\lVert(\beta_{\text{\rm init}}-\beta)_{T_{0}}\right\rVert_{2}$, and the $\ell_{1}$ (and the $\ell_{2}$) norm of the estimated coefficients on $T_{0}^{c}$, namely, $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}=\left\lVert\beta_{\text{\rm init},T_{0}^{c}}\right\rVert_{1}$ are tightly bounded with respect to $\left\lvert T_{0}\right\rvert$ for the Lasso; hence we obtain an oracle inequality on the $\ell_{2}$ loss in the sense of (11).
@@
First, by definition of $h$ and $T_{0}$, we aim to keep variables in $T_{0}$, while for variables not in $T_{0}$, we may trim these off. It is also clear by Lemma 2.3 that we cannot cut too many “significant” variables in $T_{0}$ by following the thresholding rules in our proposal; for example, for those that are $\geq\lambda\sigma\sqrt{s_{0}}$, we can cut at most a constant number of them. Now, what do we do with those in $T_{1}$? The fate of these variables is pretty much up to the choice of the threshold for a given $\beta_{\text{\rm init}}$, knowing these are the largest in magnitude in $h$ (and $\beta_{\text{\rm init}}$) outside of $T_{0}$ and hence most likely to be included in model $I$. Moreover, even if we were able to retain all variables in $T_{0}$, we will include at least some variables in $T_{1}$ when the selection set has size $\left\lvert I\right\rvert>s_{0}$; in fact, we will include nearly all variables in $T_{1}$ when the set $I$ has size close to $2s_{0}$. In particular,
@@
$\left|I\cap T^{c}_{0}\right|$ $\leq$ ${\left\lVert h_{T_{0}^{c}}\right\rVert_{1}}/{t_{0}}\leq D_{1}\lambda_{n}s_{0}/(f_{0}\lambda_{n})\leq(D_{1}/f_{0})s_{0}\quad\text{in case}\quad t_{0}=f_{0}\lambda_{n}$
@@
for some $f_{0}>0$ by (19). Before we continue, we state Lemma 3.1, which is the same (up to normalization) as Lemma 3.1 [9], to illuminate the roles of sets $T_{0},T_{1}$ in the overall bounds on $\left\lVert h\right\rVert_{2}$. We note that in their original statement, the UUP condition is assumed; a careful examination of their proof shows that it is a sufficient but not necessary condition; indeed we only need to assume that sparse eigenvalues are bounded, namely, $\Lambda_{\min}(2s_{0})>0$ and $\Lambda_{\max}(2s_{0})<\infty$. We also state Lemma 3.2.
@@
Suppose $\Lambda_{\min}(2s_{0})>0$ and $\Lambda_{\max}(2s_{0})<\infty$. Then
$\left\lVert h_{T_{01}}\right\rVert_{2}$ $\leq$ $\frac{1}{\sqrt{\Lambda_{\min}(2s_{0})}}\left\lVert Xh\right\rVert_{2}/\sqrt{n}+\frac{\theta_{s_{0},2s_{0}}}{\Lambda_{\min}(2s_{0})}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}/\sqrt{s_{0}},$ (38)
@@
${\mathcal{C}}(s_{0},k_{0}):=\left\{x\in\mathbb{R}^{p},\exists{J\subseteq[p],|J|=s_{0}}\text{ s.t.}\left\lVert x_{J^{c}}\right\rVert_{1}\leq k_{0}\left\lVert x_{J}\right\rVert_{1}\right\}.$ (40)
@@
Now we provide a proof sketch for Theorem 2.4. For $u=(u_{1},\ldots,u_{n})\in\mathbb{R}^{n}$, define the empirical norm of $u$ by $\left\lVert u\right\rVert_{n}^{2}=\left\lVert u\right\rVert_{2}^{2}/n.$ The proof of the original Theorem 5.1 in [48] draws upon techniques from a concurrent work by [36] and uses an elementary inequality for the Lasso; cf. the supplementary (D.1). Denote by $\beta_{0}:=\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$. The current proof replaces the supplementary (D.1) for the Lasso with the following updated inequality:
$\left\lVert X(\beta_{\text{\rm init}}-\beta)\right\rVert^{2}_{n}+\left\lVert X(\beta_{\text{\rm init}}-\beta_{0})\right\rVert^{2}_{n}\leq\left\lVert X(\beta-\beta_{0})\right\rVert^{2}_{n}$ (41)
$+\frac{2}{n}\epsilon^{T}X(\beta_{\text{\rm init}}-\beta_{0})+2\lambda_{n}(\left\lVert\beta_{0}\right\rVert_{1}-\lVert\beta_{\text{\rm init}}\rVert_{1}).$
See Eq(20) [11], where we set $\bar{\beta}=\beta_{0}=\beta^{\mathop{\natural}\kern 0.60275pt}({T_{0}})$ and $\bar{\delta}=\beta_{\text{\rm init}}-\beta_{0}=h$.
Denote by $\delta=\beta_{\text{\rm init}}-\beta$. Similar to the proof in [48], we have a deterministic proof on event ${\mathcal{T}}_{a}$, except that now on ${\mathcal{T}}_{a}$,
$\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}+\lambda_{n}\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\leq\left\lVert X(\beta-\beta_{0})\right\rVert^{2}_{n}+3\lambda_{n}\left\lVert h_{T_{0}}\right\rVert_{1}.$ (42)
We differentiate between three cases under event ${\mathcal{T}}_{a}$.
@@
    $\left\lVert X\delta\right\rVert^{2}_{n}+\left\lVert Xh\right\rVert^{2}_{n}\geq\left\lVert X(\beta-\beta_{0})\right\rVert_{n}^{2};$ (43)
    Then, $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\leq 3\left\lVert h_{T_{0}}\right\rVert_{1},\text{ and hence }\;\;h\in{\mathcal{C}}(s_{0},3).$
@@
    Moreover, the bounds on $\left\lVert X\delta\right\rVert_{n}$ and $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ (and $\left\lVert h\right\rVert_{1}$) follow from (45):
@@
    where $\left\lVert X(\beta-\beta_{0})\right\rVert_{n}\leq\lambda\sigma\sqrt{s_{0}\Lambda_{\max}(s-s_{0})}$. Then we have
@@
[…]
@@
    $\left\lVert h_{T_{0}}\right\rVert_{1}$ $\leq$ $\left\lVert X\beta-X\beta_{0}\right\rVert_{n}^{2}/\lambda_{n}\asymp\lambda_{n}s_{0}.$
@@
Using (45), (136), (48), and (50), we conclude that (22) holds. Although the new proof is introduced to get rid of the constant factor $2$ in front of $\left\lVert X\beta-X_{I_{0}}\beta_{I_{0}}\right\rVert_{n}=\left\lVert X(\beta-\beta_{0})\right\rVert_{n}$ in (22) in [48], we are not optimizing the constants in this paper; cf. the proof of the original Theorem 5.1 in [48]. For example, the constant $3$ in (42) can be further reduced following [11] using (125). Moreover, combining (44), (46), (49), and (51), we have the expression of (27) for $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$, namely, $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}\leq D_{1}\lambda\sigma s_{0}$. We demonstrate the tightness of these bounds in Fig. 2.
Next we bound $\left\lVert h\right\rVert_{2}$ in view of Lemma 3.1. Lemma 3.3 also provides the upper bounds on $\left\lVert h_{T_{01}}\right\rVert_{2}$, leading to the expression of $D_{0}$ as in (23). We combine the upper bounds in Lemma 3.1 and the supplementary Lemma D.2 in the proof of Lemma 3.3. We prove Lemmas 3.1 and 3.3, and Theorem 2.4 in the supplementary Sections CE, and D respectively.